OpenStudy (anonymous):

A teacher can grade 25 papers in 4 hours. How many papers can he grade in 6 hours? A. 29 B. 37.5 C. 31 D. 35

OpenStudy (anonymous):

@babygirl121314 @iGreen @DisCrazyGinger

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

25 papers in FOUR hours 25 / 2 = TWO hours 12.5 = TWO HOURS FOUR + TWO = SIX (25) + (12.5) = 37.5

OpenStudy (igreen):

25 / 4 = 6.25 6.25 \(\times\) 6 = 37.5

OpenStudy (anonymous):

25/4=? ?*6= and that's how you find it out

OpenStudy (anonymous):

37.5?

OpenStudy (anonymous):

37.5

OpenStudy (anonymous):

yes